i have this problem tat goes----
Given that u(t) = 2t² - 1,
Find the expression for u(3t) – 2u(t)

Answer by vasumathi(46)

Given that u(t) = 2t² - 1,
Find the expression for u(3t) – 2u(t)
Solution:u(t) = 2t^2-1
u(3t) =2(3t)^2-1
u(3t)=2*9t^2-1
u(3t) = 18t^2-1
2u(t)= 2(2t^2-1)= 4t^2-2
u(3t)-2u(t)= 18t^2-1-( 4t^2- 2)=18t^2-1-4t^2+2=14t^2+1
This is the final answer